What should be subtracted from 2x3-3x2y -2xy2 +3y2 to get x3 -2x2y + 3xy2 +4y2

Grade:8

1 Answers

REYAN MANZOOR
3 years ago
solution:
let the secomd number be A
according to condition
2x3-3x2y-2xy2+3y2-A=x3-2x2y+3xy2+4y2
2x3-x3-3x2y+2x2y-2xy2-3xy2+3y2-4y2=A
x3-x2y-5xy2-y2=A